# Note on two-point mean square displacement

### v1

@@ Abstract
When probe molecules of interest are embedded in a container or aggregate under stochastic motion, one needs to rely on the so-called two-point mean square displacement (MSD) measurement to extract the intrinsic mobility of the probes. We discuss two versions, based on the time series of relative vector or distance between two probes, and summarize their basic properties compared to the standard MSD. We also propose a way to extract (i) the non-Gaussianity in the displacement statistics and (ii) the motional correlation between probes from the two-point MSD. The results are presented not only for independent probes, but also for intramolecular probes within a long polymer, which could be useful in quantifying the dynamics of chromatin loci in living cell nucleus.
@@
[…]
@@
Although calculation of MSD using Eq. (1) looks quite straightforward, one may encounter situations, often in the measurement in biological samples, where it is not the case. Indeed, a complication arises when the molecule of interest is embedded in a medium (container or aggregate), which itself moves. The best example is found in the analysis of chromatin dynamics inside a cell nucleus Marshall1997 ; Mine-Hattab2012 ; Ochiai2015 ; Arai2017 ; Khanna2019 ; Yesbolatova2022 ; Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 ; we are interested in the motion of chromatin locus, but the time series of its position reflects not only the intrinsic chromatin dynamics, but also the translational and rotational motion of the nucleus, which is subjected to the random or systematic forces from cytoplasm, see Fig. 1. While such an effect of the nuclear motion might be negligible in short time scale, it becomes increasingly visible in longer time scale. For instance, in early embryos of C. elegans, the nuclear motion dominates the apparent motion of chromatin locus already at $\tau\sim 10$ seconds Yesbolatova2022 . A natural strategy here to extract the intrinsic chromatin mobility is to make use of the trajectories of two probes. The same problem is encountered in many other cases, for instance, when analyzing the motion of individual cells within multicellular aggregates. With such a situation in mind, Pönisch and Zaburdaev have recently studied the statistics of the relative distance between independently moving Brownian particles Ponisch2018 . In the chromatin problem, the polymeric effect should be taken into account, in which one may look at not only independently moving intermolecular pair but also intramolecular pair, see Fig. 2 Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 .
@@
Figure 1: Schematics illustrating the effect of (a) translational or (b) rotational motion of a container on the relative vector ${\vec{d}(t)}$ and the distance $d(t)$. Intrinsic mobility of probes are hypothetically set to zero to emphasize the extrinsic effect.
@@
Let ${\vec{r}}_{1}(t)$ and ${\vec{r}}_{2}(t)$ to represent the positions of the two probes at time $t$. The method of analysis using ${\vec{r}}_{1}(t)$ and ${\vec{r}}_{2}(t)$ to quantify the intrinsic mobility of probe may be called two-point MSD analysis. In literature, there are two variants of the two-point MSD analysis, which differ in the definition of the analyzed quantities, i.e., one may look at the time series of either a relative vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{2}(t)$ connecting two probes or its absolute value $d(t)=|{\vec{d}}(t)|$, the latter being a distance between the probes. In this note, we aim to clarify the similarity and difference between these two analyses, and their quantitative connection to the standard MSD. In Sec. I, we first define the quantities considered in two-point MSD analysis. We then briefly review the case with two independently moving probes in Sec. II. We proceed in Sec. III to the intramolecular analysis, in which two probes belong to the common polymer chain, i.e., two loci in the same chromatin. Sec. IV is devoted to discussion on the general aspect of two-point MSD. We also demonstrate that the method can be applied to quantify the non-Gaussian parameter in the displacement statistics and the motional correlation of probe pair.
@@
Figure 2: Possible variations for a pair of probes in two-point MSD measurement. (a) and (b): Pair of independent probes; while probes are simple particles in (a), tagged monomers play a role of probe in (b), each of which belongs to different polymers. (c) Instead of the second probe, one can make use of the center of mass position ${\vec{r}_{c}}(t)$ (or other fixed point) in the moving container to define a vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{c}(t)$. (d) Intramolecular probe pair as an example of mutually dependent probes.
I Quantities analyzed in two-point MSD
@@
with $\Delta d(\tau)=d(t_{0}+\tau)-d(t_{0})$. An important difference between these quantities is found by observing that MSCV excludes only the translational, but not the rotational motion of the container, see Fig. 1. Thus, the MSCD is more suitable when the rotational motion of the container is non-negligible.
II Independently moving probes
II.1 Relative vector analysis
@@
Figure 3: Two-point MSD analysis for a pair of independent probes. (a) Time evolution of the probability density function $P_{d}(d,t|d_{0})$ of the inter-probe distance $d$ starting from $d(0)=d_{0}$. (b) Conditional ${\rm MSCD}(\tau|d_{0})$ as a function of lag time $\tau$ for two different initial separation $d_{0}$. (c)${\rm MSCV}(\tau)$ and ${\rm MSCD}(\tau)$ as a function of lag time $\tau$ compared to ${\rm MSD}(\tau)$. In (b) and (c), the length is measured in unit of $a$ (probe diameter) and the time is measured in unit of $a^{2}/2D^{*}$.
II.2 Distance analysis
@@
The time evolution of $P_{d}(d,\tau|d_{0})$ is shown in Fig. 3 (a), where the width of the distribution grows as $\sim\tau^{1/2}$. From Eq. (8), we calculate the MSCD under the condition $d(0)=d_{0}$;
@@
$q_{>}=\int_{\sqrt{4D^{*}\tau}}^{\infty}\ p(d_{0})\ {\mathrm{d}}d_{0},\ q_{<}=\int_{0}^{\sqrt{4D^{*}\tau}}\ p(d_{0})\ {\mathrm{d}}d_{0}$ (15)
For a pair of free probes, the relative vector ${\vec{d}}$ is uniformly sampled. If we tentatively introduce the maximum $d_{m}$ for the allowed $d_{0}$, then
@@
For any given finite $\tau$, $q_{>}\rightarrow 1$ and $q_{<}\rightarrow 0$ in large enough space $d_{m}\rightarrow\infty$. We thus find
@@
for two independently moving probes, see Fig. 2 (a). If two probes are identical, we find ${\rm MSCD}(\tau)=4D\tau=2/3{\rm MSD}(\tau)$.
II.3 Motion analysis in center-of-mass frame
Instead of tracking two probes, one can track the position of a single probe ${\vec{r}}_{1}(t)$ with respect to a particular fixed point in the moving container. A convenient choice for the fixed point would be the center of mass, whose position ${\vec{r}}_{c}(t)$ dynamically changes with the container’s translation Ochiai2015 . Similar to the two-probes tracking case, one can define the vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{c}(t)$, and its magnitude $d(t)=|{\vec{d}}(t)|$, see Fig. 2 (c). In order to avoid a possible confusion, let us attach a subscript “c” as ${\rm MSCV_{c}}$ and ${\rm MSCD_{c}}$ to indicate the quantities of interest in the center-of-mass frame analysis. Since the diffusivity of the fixed point (second probe) is obviously zero, the relative diffusion coefficient in this case is simply $D^{*}=D_{1}$. Thus, we find from Eq. (7)
${\rm MSCV_{c}}(\tau)=6D_{1}\tau={\rm MSD}(\tau)=\frac{1}{2}{\rm MSCV}(\tau)$ (18)
@@ II.3 Motion analysis in center-of-mass frame
${\rm MSCD_{c}}(\tau)=2D_{1}\tau=\frac{1}{3}{\rm MSD}(\tau)=\frac{1}{2}{\rm MSCD}(\tau)$ (19)
III Intramolecular probes
@@
Model—. Our calculation in this section is based on the Rouse model. The polymer consists of $N$ monomers connected through spring in series. The position of $n$-th monome ${\vec{r}}(n,t)=(x(n,t),y(n,t),z(n,t))$ evolves according to the Rouse equation of motion, which takes in continuum limit the following form
@@
where the friction coefficient $\gamma$ and the spring constant $k$ defines the monomeric time scale $\tau_{0}=\gamma/k$. The last term ${\vec{\xi}}(n,t)$ is uncorrelated Gaussian white noise acting on $n$-th monomer with $\langle{\vec{\xi}}(n,t)\rangle={\vec{0}}$ and $\langle{\vec{\xi}}(n,t){\vec{\xi}}(m,s)\rangle=2\gamma k_{B}T{\mathcal{I}}\delta(n-m)\delta(t-s)$, which represents the effect of thermal noise, with $k_{B}T$ being the thermal energy and $3\times 3$ identity matrix ${\mathcal{I}}$.
@@
[…]
@@
where in the final step, we focus on the stationary state by letting $t_{0}\rightarrow-\infty$, which implies $x(n,t+\tau)-x(n,t)=\delta x(n,t+\tau)-\delta x(n,t)$. This leads to the well-known subdiffusion scaling of MSD of each probe (with exponent $0.5$ for a Rouse model) DoiEdwards ; Panja2010 ; Saito2015 ;
@@
$=$ $\frac{2a^{2}}{\sqrt{\pi}}\sqrt{\frac{\tau}{\tau_{0}}},$
where we introduce the mean square bond length $a^{2}$ through equipartition theorem $k=3k_{B}T/a^{2}$ and a factor $3$ reflects the dimensionality (the number of components) and the isotropy of the space.
Figure 4: Two-point MSD analysis for a pair of intramolecular probes. (a) and (b) Time evolution of the probability density function $P_{d}(d,t|d_{0})$ of the inter-probe distance $d$ starting from (a) $d(0)=d_{0}=0.5\sqrt{ma^{2}}$ and (b) $d(0)=d_{0}=1.5\sqrt{ma^{2}}$ Dashed curves represent the equilibrium distribution. (c) Conditional ${\rm MSCD}(\tau|d_{0})$ as a function of scaled lag time $\tau/\tau_{m}$ for three different initial separation $d_{0}$. (d) ${\rm MSCD}(\tau)$ as a function of lag time $\tau$ for three different inter-probe separation $m$. (e)${\rm MSCV}(\tau)$ and ${\rm MSCD}(\tau)$ as a function of scaled lag time $\tau/\tau_{m}$ compared to ${\rm MSD}(\tau)$.
@@
III.1 Relative vector analysis

### v2

@@ Abstract
When probe molecules of interest are embedded in a substrate or aggregate under stochastic motion, one needs to rely on the so-called two-point mean square displacement (MSD) measurement to extract the intrinsic mobility of the probes. We discuss two versions, based on the time series of relative vector or distance between two probes, and summarize their basic properties compared to the standard MSD. We also propose a way to extract (i) the non-Gaussianity in the displacement statistics and (ii) the motional correlation between probes from the two-point MSD. The results are presented not only for independent probes, but also for intramolecular probes within a long polymer, which could be useful in quantifying the dynamics of chromatin loci in living cell nucleus.
I Introduction
@@
where $\Delta{\vec{r}}(\tau)={\vec{r}}(t_{0}+\tau)-{\vec{r}}(t_{0})$. MSD thus measures the typical displacement squared during the time interval $\tau$. Note that MSD generally depends on the time origin $t_{0}$. In the present note, we assume stationary state, where time translational invariance implies no dependence of statistical quantities on $t_{0}$, and $\langle\cdots\rangle$ represents the average over trajectories and over the time origin $t_{0}$.
Although calculation of MSD using Eq. (1) looks quite straightforward, one may encounter situations, often in the measurement in biological samples, where it is not the case. Indeed, a complication arises when the molecule of interest is embedded in a medium or an aggregate, hereafter called substrate, which itself moves. In such a case, the observed displacement $\Delta{\vec{r}}^{(o)}(\tau)$ is a superposition of two contributions;
$\Delta{\vec{r}}^{(o)}(\tau)=\Delta{\vec{r}}(\tau)+\Delta{\vec{r}}^{(ex)}(\tau)$ (2)
@@
where $\Delta{\vec{r}}(\tau)$ represents the displacement due to the intrinsic dynamics of probe, while $\Delta{\vec{r}}^{(ex)}(\tau)$ stands for the extrinsic contribution arising from the substrate motion. If one uses $\Delta{\vec{r}}^{(o)}(\tau)$ in Eq. (1), the resultant observed MSD, which we denote as ${\rm MSD}^{(o)}(\tau)$, would be different from ${\rm MSD}(\tau)$, thus fails to capture the intrinsic mobility of the probe. The best example is found in the analysis of chromatin dynamics inside a cell nucleus Marshall1997 ; Mine-Hattab2012 ; Ochiai2015 ; Arai2017 ; Khanna2019 ; Yesbolatova2022 ; Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 ; we are interested in the motion of chromatin locus, but the time series of its position reflects not only the intrinsic chromatin dynamics, but also the translational and rotational motion of the nucleus, which is subjected to the random or systematic forces from cytoplasm, see Fig. 1. While such an effect of the nuclear motion might be negligible in short time scale, it becomes increasingly visible in longer time scale. For instance, in early embryos of C. elegans, the nuclear motion dominates the apparent motion of chromatin locus already at $\tau\sim 10$ seconds Yesbolatova2022 . The relevance of nuclear motion has also been reported in chromatin dynamics in mouse embryonic stem cells Oliveira2021 . A natural strategy here to extract the intrinsic chromatin mobility is to make use of the trajectories of two probes. The same problem is encountered in many other cases, for instance, when analyzing the motion of individual cells within multicellular aggregates. With such a situation in mind, Pönisch and Zaburdaev have recently studied the statistics of the relative distance between independently moving Brownian particles Ponisch2018 . In the chromatin problem, the polymeric effect should be taken into account, in which one may look at not only independently moving intermolecular pair but also intramolecular pair, see Fig. 2 Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 .
@@
Figure 1: Schematics illustrating the effect of (a) translational or (b) rotational motion of a substrate on the relative vector $\vec{d}(t)$} and the distance $d(t)$. Intrinsic mobility of probes are hypothetically set to zero to emphasize the extrinsic effect.
@@
Let ${\vec{r}}_{1}(t)$ and ${\vec{r}}_{2}(t)$ to represent the positions of the two probes at time $t$. The method of analysis using ${\vec{r}}_{1}(t)$ and ${\vec{r}}_{2}(t)$ to quantify the intrinsic mobility of probe may be called two-point MSD analysis. We note that the term two-point MSD has so far been used in a loose sense without a proper definition. In literature, there are two variants of the two-point MSD analysis, which differ in the definition of the analyzed quantities, i.e., one may look at the time series of either a relative vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{2}(t)$ connecting two probes or its absolute value $d(t)=|{\vec{d}}(t)|$, the latter being a distance between the probes. In this note, we aim to clarify the similarity and difference between these two analyses, and their quantitative connection to the standard MSD. In Sec. II, we first define the quantities considered in two-point MSD analysis, and present a short summary of the paper. We then proceed to the analysis of simple solvable models, first the case with two independently moving probes in Sec. III and then two probes belonging to the common polymer, i.e., two loci in the same chromatin in Sec. IV. Sec. V is devoted to discussion on the general aspect of two-point MSD. We also demonstrate that the method can be applied to quantify the non-Gaussian parameter in the displacement statistics and the motional correlation of probe pair.
@@
Figure 2: Possible variations for a pair of probes in two-point MSD measurement. (a) and (b): Pair of independent probes; while probes are simple particles in (a), tagged monomers play a role of probe in (b), each of which belongs to different polymers. (c) Instead of the second probe, one can make use of the center of mass position ${\vec{r}_{c}}(t)$ (or other fixed point) in the moving substrate to define a vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{c}(t)$. (d) Intramolecular probe pair as an example of mutually dependent probes.
II Short summary
In this section, we first define the central quantities analyzed in two-point MSD analysis, and present a short summary of the present paper. More detailed analyses and discussions can be found in subsequent sections. Readers who are interested only in the general aspects of two-point MSD analysis may skip Secs. III and IV, where detailed analyses on simple solvable models are presented.
II.1 Quantities analyzed in two-point MSD
@@
with $\Delta d(\tau)=d(t_{0}+\tau)-d(t_{0})$. One can rewrite the change in the vector as $\Delta{\vec{d}}(\tau)=\Delta{\vec{r}}_{1}(\tau)-\Delta{\vec{r}}_{2}(\tau)$, which indicates that $\Delta{\vec{d}}(\tau)$ may be also influenced by the substrate motion. Therefore, similarly to ${\rm MSD}^{(o)}(\tau)$, we also define ${\rm MSCV}^{(o)}(\tau)$ and ${\rm MSCD}^{(o)}(\tau)$, which are obtained from the observed displacement $\Delta{\vec{r}}_{i}^{(o)}(\tau)$ for $i=1,2$.
An important difference between these quantities is found by observing that MSCV excludes only the translational, but not the rotational motion of the substrate, i.e., ${\rm MSCV}^{(o)}(\tau)={\rm MSCV}(\tau)$ only if the rotational motion of the substrate is negligible see Fig. 1. On the other hand, the equality ${\rm MSCD}^{(o)}(\tau)={\rm MSCD}(\tau)$ always holds irrespective of the type of substrate motion.
Figure 3: (a)MSD, (b)MSCV and (c)MSCD of probe (or probe pair) performing fractional Brownian motion with MSD exponent $\alpha=0.5$ under the influence of stochastic substrate motion; (left) translational diffusion and (right) rotational diffusion. In each case, both the apparent and the intrinsic quantities are plotted. The former is calculated from the observed displacement $\Delta{\vec{r}}^{(o)}(\tau)$ and marked by the superscript $(o)$, see the main text for details. Length is measured in unit of probe size, and the unit of time is set by the corresponding time-scale, i.e., the time interval for the probe to make displacement over its own size, see Appendix VI.2 for details of the model.
As a demonstration, we have carried out a simple simulation for probe particles in a moving substrate (see Appendix VI.2 for simulation details). As a model of the intrinsic probe motion, we have adopted the sub-diffusional fractional Brownian motion, which mimics the motion of a tagged monomer in polymer. In Fig. 3 (a), we show results of MSD calculated from $\Delta{\vec{r}}(\tau)$ or $\Delta{\vec{r}}^{(o)}(\tau)$, where $\Delta{\vec{r}}^{(o)}(\tau)$ includes the contribution from the substrate translational (left) or rotational (right) diffusive motion. In both cases, while the substrate motion is negligible in a short time scale, it eventually dominates over the intrinsic mobility in a longer time scale, leading to a spurious higher MSD exponent. It is expected that the more vigorous the substrate motion is, the shorter this crossover time scale (see Appendix VI.2 for the estimation of the time scale). Figs. 3 (b) and (c) show the same analysis for MSCV and MSCD, respectively. Clearly, while MSCV is capable of eliminating the substrate translational motion only, MSCD succeeds in quantifying the intrinsic probe mobility embedded in a substrate undergoing translational and rotational motions.
II.2 Relation between MSD, MSCV and MSCD
Given the uncontrollable dependence of ${\rm MSD}^{(o)}$ on the substrate motion, and the ability of its partial or complete removal in two-point MSD analysis, it is useful to find a formula to obtain MSD from MSCD or MSCV. If probes are uncorrelated in their motion, such a formula is given by
${\rm MSD}(\tau)=\frac{1}{2}{\rm MSCV}(\tau)=\frac{3}{2}{\rm MSCD}(\tau)\$(5)
where the mobilities of a pair of two probes are assumed to be equal. One can provide simple and intuitive explanations for the relation. First, ${\rm MSCV}={\rm MSD}_{1}+{\rm MSD}_{2}$, where ${\rm MSD}_{i}$ is the MSD of $i$-th probe, as long as two probes are uncorrelated. It follows directly from the fact that variances of uncorrelated random variables add. Second, ${\rm MSCD}={\rm MSCV}/3=({\rm MSD}_{1}+{\rm MSD}_{2})/3$ follows from the fact that changes in the vector ${\vec{d}}={\vec{r}}_{1}-{\vec{r}}_{2}$ can be decomposed into radial (1 dimension) and tangential (3-1=2 dimensions), where MSCD is sensitive only to the radial component in short time scale. The precise meaning of the short time scale, hence the applicability of Eq. (5) for MSCD will be clarified in the following sections using simple concrete models in Secs. III and IV. A more general discussion is given in Sec. V.
III Independently moving probes
In this section, we present a detailed two-point MSD analysis for the simplest model situation, where probe particles perform independent Brownian motion. The case of intramolecular probe pair within a polymer will be treated in Sec. IV.
III.1 Relative vector analysis
@@
[…]
III.2 Distance analysis
@@
The time evolution of $P_{d}(d,\tau|d_{0})$ is shown in Fig. 4 (a), where the width of the distribution grows as $\sim\tau^{1/2}$. From Eq. (LABEL:P_d_indep), we calculate the MSCD under the condition $d(0)=d_{0}$;
@@
$q_{>}=\int_{\sqrt{4D^{*}\tau}}^{d_{m}}\ p(d_{0})\ {\mathrm{d}}d_{0}$ (17)
and $q_{<}=1-q_{>}$, where we introduce the maximum $d_{m}$ for the allowed $d_{0}$. For a pair of free probes, the relative vector $\vec{d}$} is uniformly sampled, thus
@@
At short time scale $\tau\ll d_{m}^{2}/4D^{*}$, one then finds $q_{>}\rightarrow 1$, hence,
@@
for two independently moving probes, see Fig. 2 (a). In this short time scale regime, the particles move much less than their average initial separation. This regime grows with $d_{m}$ and extends indefinitely if the particles are on average infinitely far apart. If two probes are identical, we find ${\rm MSCD}(\tau)=4D\tau=2/3{\rm MSD}(\tau)$.
III.3 Motion analysis in center-of-mass frame
Instead of tracking two probes, one can track the position of a single probe ${\vec{r}}_{1}(t)$ with respect to a particular fixed point in the moving substrate. A convenient choice for the fixed point would be the center of mass, whose position ${\vec{r}}_{c}(t)$ dynamically changes with the substrate translation Ochiai2015 . Similar to the two-probes tracking case, one can define the vector ${\vec{d}}(t)={\vec{r}}_{1}(t)-{\vec{r}}_{c}(t)$, and its magnitude $d(t)=|{\vec{d}}(t)|$, see Fig. 2 (c). In order to avoid a possible confusion, let us attach a subscript “c” as $\rm MSCV_{c}$} and $\rm MSCD_{c}$} to indicate the quantities of interest in the center-of-mass frame analysis. Since the diffusivity of the fixed point (second probe) is obviously zero, the relative diffusion coefficient in this case is simply $D^{*}=D_{1}$. Thus, we find from Eq. (9)
${\rm MSCV_{c}}(\tau)=6D_{1}\tau={\rm MSD}(\tau)=\frac{1}{2}{\rm MSCV}(\tau),$ (20)
@@ III.3 Motion analysis in center-of-mass frame
${\rm MSCD_{c}}(\tau)=2D_{1}\tau=\frac{1}{3}{\rm MSD}(\tau)=\frac{1}{2}{\rm MSCD}(\tau).$ (21)
Note that Eq. (21) applies in the short time scale regime, see discussion around Eq. (19).
IV Intramolecular probes
@@
Model—. Our calculation in this section is based on the Rouse model. The polymer consists of $N$ monomers connected through spring in series. The position of $n$-th monome ${\vec{r}}(n,t)=(x(n,t),y(n,t),z(n,t))$ evolves according to the Rouse equation of motion, which takes the following form
@@
where the friction coefficient $\gamma$ and the spring constant $k$ defines the monomeric time scale $\tau_{0}=\gamma/k$. The last term ${\vec{\xi}}(n,t)$ is uncorrelated Gaussian white noise acting on $n$-th monomer with $\langle{\vec{\xi}}(n,t)\rangle={\vec{0}}$ and $\langle{\vec{\xi}}(n,t){\vec{\xi}}(m,s)\rangle=2\gamma k_{B}T{\mathcal{I}}\delta(n-m)\delta(t-s)$, which represents the effect of thermal noise, with $k_{B}T$ being the thermal energy and $3\times 3$ identity matrix $\mathcal{I}$}. Note that the above continuum approximation (22) is valid on length scale down to the monomer size $a$ with $k=3k_{B}T/a^{2}$.
@@
${\vec{r}}(n,t)$ $=$ $\frac{1}{\gamma}\int dn^{\prime}\int_{t_{0}}^{t}dt^{\prime}\ G(n-n^{\prime},t-t^{\prime}){\vec{\xi}}(n^{\prime},t^{\prime})$ (23)
$+$ $\int dn^{\prime}\ G(n-n^{\prime},t-t_{0}){\vec{r}}(n^{\prime},t_{0})$
@@
where in the final step, we focus on the stationary state by letting $t_{0}\rightarrow$-$\infty$, in which the time translational invariance indicates $\langle x(n,t+\tau)\rangle=\langle x(n,t)\rangle$, hence, $x(n,t+\tau)-x(n,t)=\delta x(n,t+\tau)-\delta x(n,t)$. This leads to the well-known subdiffusion scaling of MSD of each probe (with exponent $0.5$ for a Rouse model) DoiEdwards ; Panja2010 ; Saito2015 ;
@@
$=$ $\frac{2a^{2}}{\sqrt{\pi}}\sqrt{\frac{\tau}{\tau_{0}}},$ (29)
where a factor $3$ reflects the dimensionality (the number of components) and the isotropy of the space, and we express the final expression in terms of monomer size $a$ and the corresponding time scale $\tau_{0}$.
[…]
@@
IV.1 Relative vector analysis